Briefly discuss a strategy a firm operating in a perfectly competitive market must take to sustain its profit outlook.                                                                                           [3 Marks]
A firm operating in a competitive market has the following cost function C = 10 + 2Q2. The market demand is Q = 40 – 2P and equilibrium price of output is K20.
Calculate the output the firm must to maximize its profits.                        [3 Marks]
Find the price the firm must charge to maximize profits.                            [2 Marks]
Answer #1

Q1. In perfectly competitive market, the marginal revenue is equal to the price since the demand schedule is horizontal. The firm can maximize its profit at the point where MR=P=MC. Apart from it, having a more efficient production technology than rivals will help in cost-cutting which will ensure higher profits.

Q2. C = 10 + 2Q2. So, MC = 2

Q = 40-2P

So, P = 20-0.5Q

Therefore, MR = 20-Q

MR=MC [Profit maximization condition]

This implies 20 - Q = 2

So, Q = 18

and P = 11